How do you determine the solutions of quadratic equations? How about rational algebraic equations transformable to quadratic equations?
Answer by Theo(13342)

you need to put the quadratic equation into standard form.

the standard form is ax^2 + bx + c = 0

a is the coefficient of the x^2 term.

b is the coefficient of the x term.

c is the constant term.

once you have it in that form, you can solve it using the quadratic formula.

that formula is:



        -b plus or minus square root of (b^2 - 4ac)
x = ---------------------------------------------------
                             2a



here's a reference.

https://www.mathsisfun.com/algebra/quadratic-equation.html

here's a bunch of references that will tell you a lot about how to solve quadratic equations.

https://www.google.com/search?q=solving+quafratic+equations+purplemath&rlz=1C1GCEA_enUS874US874&oq=solving+quafratic+equations+purplemath&aqs=chrome..69i57j0i22i30i457j0i10i22i30.6751j0j7&sourceid=chrome&ie=UTF-8

the definition of a quadratic equation is that the highest order of degree is 2.

if the equation is in one variable only, then the degree of the equation is the term that has the largest exponent.

for example:

y = x^2 + x + 3
y = x^4 + x^3 + x^2 + 2
y = x^7 + 15

the degree of the first equation is 2 because the term with the highest exponents is 2.
the degree of the second equation is 4 because the term with the highest exponent is 4.
the degree of the third equation is 7 because the term with the highest exponent is 7.

a quadratic equation is has a degree of 2.

y = x^2 is a quadratic
y = x^2 + 15x - 3 is a quadratic
y = x^2 + 5 is a quadratic.

y = x + 3 is not
y = x^4 - 3 is not
etc.

the quadratic formula is your last resort.
if you can't solve it by any other method, then use the quadratic formula.